How much more does $1000 earn in eight years, compounded daily at 5% than $1000 over eight years at 5% compounded semiannually
Answer by Boreal(15235)

Daily: A=5000(1+(.05/365)^8*365=$7458.92, rounding at the end. We are raising to the 2920 power.
Semi-annual: A=5000(1+0.025)^16=$7422.53
That is a difference of $36.39